OpenStudy (boldjon):

A personal trainer designs exercise plans based on a combination of strength-training and aerobic exercise. A beginner plan has 15 minutes per session of strength-training and 20 minutes per session of aerobic exercise for a total of 90 minutes of exercise in a week. An advanced plan has 20 minutes per session of strength-training and 30 minutes of aerobic exercise for a total of 130 minutes of exercise in a week.

OpenStudy (boldjon):

Which statement describes when the plans are based on the same number of aerobic exercise sessions? a.Each plan utilizes a combination of 2 strength-training sessions and 2 aerobic exercise sessions per week. b. Each plan utilizes a combination of 2 strength-training sessions and 3 aerobic exercise sessions per week. c. Each plan utilizes a combination of 3 strength-training sessions and 2 aerobic exercise sessions per week. d. Each plan utilizes a combination of 3 strength-training sessions and 3 aerobic exercise sessions per week.

OpenStudy (trojanpoem):

For a beginner like you : Strength training = x Aerobics = y 15x + 20y = 90 For Ultimate warriors (me): 20x + 30y= 130 Test with the given values: 15 * 2 + 20 * 3 = 90 = R.H.S 20 * 2 + 30 * 3 = 130 = R.H.S

OpenStudy (boldjon):

so...b?

OpenStudy (alivejeremy):

yea

OpenStudy (boldjon):

thanks man

OpenStudy (alivejeremy):

ok kool thts only cause u help me this moring don't fill ur self bro

OpenStudy (alivejeremy):

and fan me to

OpenStudy (alivejeremy):

peace

OpenStudy (boldjon):

just did